# Unified Medical Image Pre-training in Language-Guided Common Semantic Space

### v1

@@ 3.1 Attentive slice selection
where $l$ denotes the layer index; $h$ denotes the attention head index; $\mathbf{f}_{lh}^{q}(CLS)$ denotes the query embedding of the $\left[CLS\right]$ token at Layer $l$ and Head $h$; $\mathbf{f}_{lh}^{k}(P)$ denotes the key embedding of Layer $l$ and Head $h$ for an 3D image token at location $P$; $C$ is the number of channels for the query and key embedding.
The important slices located strategy is based on the token-level score. Each token in the original CT volume represents a small voxel. By aggregating scores based on the slice dimension, we can calculate the total score for each group of slices:
$s_{i}=\frac{1}{N}\sum_{j=1}^{N}s^{P}_{ij},$ (2)
where $s_{i}$ is the attentive score for the $i$-th slice, $s^{P}_{ij}$ is the token-level attentive score for the $j$-th voxel in $i$-th slice, $N$ represents the total number of voxels included in a slice. After aggregating the attentive scores, we can obtain text relevance scores for each 2D slice. We then choose the top $k$ slices to establish a connection with the 3D data and the report, allowing us to learn a shared feature space.
Figure 4: Attentive slice selection from 3D volume. Generally, the slice is selected according to the attention weights of the $\left[CLS\right]$ token attending to other tokens, and the $\left[CLS\right]$ token is also guided by language in the report. We compute the average attention weights within each sliced area, and then select the top K slices with the highest scores.
3.2 Cross-Dimensional Medical Visual-Language Pretraining.
We use CLIP Radford et al. (2021) loss for cross-modal pre-training of 2D and 3D medical images and their corresponding reports. CLIP is a powerful tool that enables the alignment of features from two modalities after large-scale contrastive learning. For 2D X-ray training, we directly use $T_{2D}$ and $E_{v}$ for feature extraction, obtaining the global image feature $\left[CLS\right]$ token, and then aligning it with the language encoder $E_{l}$’s $\left[CLS\right]$ token. For the training of 3D CT scan data, the 2D slices within it also carry the content of the same radiology report, so we select attentive 2D slices according to the method in Section 3.1 as joint input. Through this approach, we bring the 2D slice features and 3D features closer to the same language encoder’s features, using radiology reports as a medium to form cross-dimensional interactions.
A highlight of our work is the use of attentive slices selection to ensure that the selected 2D slices are sufficiently representative. Only in this way can these 2D slices carry the supervision information from the report and, together with the 3D features, construct a joint feature space. If we were to use random selection, it would be easy to cause mismatches between the visual and textual information, and the noise generated would make the model’s understanding on 2D data very confusing. Once the common coordinates from the report are no longer accurate, it would not be possible to effectively form a cross-dimensional information bridge.
3.3 Enhancing Dimensional Interactions via Self-distillation
In Section 3.1, we introduced the method for selecting 2D slices that can share the same report. Then, in Section 3.2, we aligned them across dimensions using text as shared coordinates for visual-textual training. In fact, apart from using text as a medium, the projected representative 2D slice features and 3D features with global information also possess strong correlations. We aim to construct an auxiliary task to directly leverage this correlation, further enhancing the cross-dimensional communication.
We adopted a simple and straightforward auxiliary task design: mask and recovery. We chose to use the self-distillation method for implementation Yang et al. (2023); Zhou et al. (2021), due to its simplicity and effectiveness. During the training process, we mask a certain proportion of both 2D and 3D tokens in the online encoder, while keeping the complete input in the EMA encoder. Therefore, this non-trivial task requires predicting the EMA encoder’s features directly from the online encoder, as there is a significant amount of missing information. For both 2D and 3D recovery tasks, the model has to learn the correlation with the other modality to obtain more reference information, thus directly strengthening the interaction between 2D and 3D features within the network.
Similarly, during the token masking phase, we also employed the attentive selection design. While passing through the EMA encoder, we calculated the patch scores as described in Equation 1, and retained the portion with the highest scores. This approach minimizes the disruption of effective lesion structures, thereby avoiding ambiguity and making the cross-modal interaction more meaningful.
During the feature distillation process, we utilized the head and loss from BYOL Grill et al. (2020). We applied this loss to both the global $\left[CLS\right]$ tokens and all local patch tokens in the output 2D and 3D features, thereby enabling interaction at different granularities to enhance feature robustness.
@@ 3.3 Enhancing Dimensional Interactions via Self-distillation
DSVE Engilberge et al. (2018) 50.1 51.0 51.5 49.7 52.1 57.8 - - -
VSE++ Faghri et al. (2017) 50.3 51.2 52.4 49.4 57.2 57.9 - - -
GLoRIA Huang et al. (2021) 86.6 87.8 88.1 86.1 88.0 88.6 67.3 77.8 89.0
@@
Caption-Transformer Cornia et al. (2020) 77.2 82.6 83.9 - - - - - -
Caption-LSTM Xu et al. (2015) 85.2 85.3 86.2 - - - - - -
Contrastive-Binary Tan & Bansal (2019) 84.5 85.6 85.8 - - - - - -
ConVIRT Zhang et al. (2022) 85.9 86.8 87.3 77.4 80.1 81.3 72.5 82.5 92.0
GLoRIA-MIMIC Huang et al. (2021) 87.1 88.7 88.0 87.0 89.4 90.2 66.5 80.5 88.8
MGCA (ResNet-50) Wang et al. (2022) 87.6 88.0 88.2 88.6 89.1 89.9 72.0 83.5 90.5
MGCA (ViT-B/16) Wang et al. (2022) 88.8 89.1 89.7 89.1 89.9 90.8 74.8 84.8 92.3
UniMedI (Ours, ViT-B/16) 89.4 89.7 90.5 90.0 90.4 91.5 80.3 92.4 94.6
Table 1: Linear classification results on CheXpert, RSNA and COVIDx with 1%, 10%, 100% training data. Area under ROC curve (AUROC [%]) are reported for CheXpert and RSNA dataset, and accuracy (ACC [%]) is reported for COVIDx dataset. The best results are highlighted in boldface.
4 Experiments
We build our universal medical framework UniMedI and pre-train on the two medical vision-report datasets with different modalities including 2D X-rays and 3D CT scans. Furthermore, extensive experiments on multiple cross-modal downstream dataset from diverse tasks are conducted to verify the effectiveness of the multi-modal vision representations. In the following subsections, we first present the pre-training experiments settings in Section 4.1 and two main downstream tasks in Section 4.2. In addition, we compare the performance of our proposed approach with the state-of-the-art vision-language processing methods in Section 4.3. Finally, we perform plenty of ablation experiments on multi-modal downstream tasks and visualization to show the validity of each module of our framework.
4.1 Pre-Training Setup
Dataset  We pre-train our UniMedI framework on the JPG version of 2D X-rays dataset MIMIC-CXR 2.0.0 Johnson et al. (2019) and the MINC version of 3D CT scans dataste BIMCV de la Iglesia Vayá et al. (2021). As the downstream 2D datasets only encompass frontal-view chest images, we remove the lateral-view images to preprocess the 2D dataset MIMIC-CXR 2.0.0. Similarly, as the downstream 3D datasets only encompass frontal-view chest images, we remove the lateral-view images to preprocess the 3D dataset BIMCV. For the processing of text reports, we remove the reports which are less than 3 tokens for 2D and 3D datasets following Wang et al. (2022).
Implementation Details  Following Gloria Huang et al. (2021), we utilize ViT-B/16 Dosovitskiy et al. (2020) as the vision encoder to extract representations in the common feature space for 2D and 3D visual data. We use BioClinicalBERT Alsentzer et al. (2019) as the text encoder to obtain the report embeddings.
Table 2: Linear classification results on CC-CCII with 1%, 10%, 100% training data. Accuracy are reported for the dataset. $\ast$ denotes the input size $16\times 96\times 96$. Others is $32\times 128\times 128$. The best results are highlighted in boldface.
@@ 4.1 Pre-Training Setup
Table 2: Linear classification results on CC-CCII with 1%, 10%, 100% training data. Accuracy are reported for the dataset. $\ast$ denotes the input size $16\times 96\times 96$. Others is $32\times 128\times 128$. The best results are highlighted in boldface.
Table 3: Classification results on CC-CCII, RICORD with full training data. ACC [%] is reported for CC-CCII and AUC [%] is reported for LUNA2016-v2. The best results are highlighted in boldface.
Table 4: Ablation study of training mode on linear classification (2D dataset CheXpert, RSNA and 3D dataset CC-CCII) settings. We report Area under ROC curve (AUROC [%]) on CheXpert and RSNA datasets, and (Acc [%]) on CC-CCII dataset. Best results of each setting are in boldface.
Training tasks CheXpert (AUC) RSNA (AUC) CC-CCII (Acc)
2D 3D $1\%$ $10\%$ $100\%$ $1\%$ $10\%$ $100\%$ $1\%$ $10\%$ $100\%$
87.1 88.0 88.4 88.7 89.5 90.3 - - -
- - - - - - 55.6 71.7 76.4
87.4 88.1 88.5 88.9 89.3 90.6 72.4 80.0 86.2
4.2 Downstream Tasks and Experimental Setup
Medical Classification  We conduct medical image classification on three representative datasets: (1) CheXpert Irvin et al. (2019), which contains 191,229 frontal-view chest radiographs. The task is to classify each image into 5 individual binary labels: atelectasis, cardiomegaly, consolidation, edema, and pleural effusion. Following Zhang et al. (2022); Huang et al. (2021), we hold out the expert-labeled validation set as test data and randomly select 5,000 radiographs from training data for validation. (2) RSNA Pneumonia Shih et al. (2019). We use the stage 2 version, which contains around 29,700 frontal view chest radiographs. The task is a binary classification, i.e., classifying each chest image into normal or pneumothorax positive. Following Huang et al. (2021), we manually split the dataset into training, validation, and test set with 70%/15%/15% ratio. (3) COVIDx Wang et al. (2020), which contains over 30k CXR images from a multinational cohort of over 16,600 patients. This dataset contains 16,490 positive COVID-19 images from over 2,800 patients. We use the latest version 6 of this dataset. The task is a three-class classification, i.e., classifying each radiograph into COVID-19, non-COVID pneumonia or normal. We use the original validation dataset as test data and manually split 10% of original training set for validation.
Table 5: Ablation study of our framework on linear classification (2D dataset CheXpert, RSNA and 3D dataset CC-CCII) settings. We report Area under ROC curve (AUROC [%]) on CheXpert and RSNA datasets, and (Acc [%]) on CC-CCII dataset. $VL$ represents the default experiment setting include image-text contrastive loss $L_{vl}$ with random slices selection. $FD$ will include $L_{icl}$ and $L_{pcl}$ loss to execute self feature distillation. $Attn$ will use attentive slices selection instead of random. Best results of each setting are in boldface.
Training tasks CheXpert (AUC) RSNA (AUC) CC-CCII (Acc)
$VL$ $FD$ $Attn$ $1\%$ $10\%$ $100\%$ $1\%$ $10\%$ $100\%$ $1\%$ $10\%$ $100\%$
87.4 88.1 88.5 88.9 89.3 90.6 72.4 80.0 86.2
89.0 89.3 90.1 89.5 90.1 91.2 74.6 80.9 86.7
89.4 89.7 90.5 90.0 90.4 91.5 75.6 84.8 89.4
We conduct medical volume classification on two representative datasets: (1) CC-CCII Zhang et al. (2020) and LUNA 16 Setio et al. (2017). More details about the 3D datasets are in Appendix.
we use the Linear Classification setting to evaluate the representative ability of our universal vision-language pre-training framework. Apart from this, we also apply Classification to evaluate UniMedI for 3D data. Linear Classification freezes the pre-trained ViT vision encoder and only training a randomly initialized linear classification head for the downstream classification task with 1%, 10%, and 100% training data on each classification dataset.
Medical Semantic Segmentation  We conduct experiments to evaluate the performance of our framework for medical semantic segmentation on RSNA and BCV datasets: (1) RSNA Pneumonia Shih et al. (2019), contains 29700 frontal view radiograph. The task is to predict bounding boxes indicating evidence of pneumonia. We randomly split the original training set into 16,010/5,337/5,337 for training/validation/testing. We convert object detection ground truths into masks for semantic segmentation. (2) BCV Landman et al. (2015), which consists of 50 CT scans and is divided into 24/26 for training/testing following Xie et al. (2022).
4.3 Result
4.3.1 Results on Medical Classification
@@
2D Medical Image Classification  Table 1 reports the results of Linear Classification on three 2D medical image classification datasets (CheXpert, RSNA and COVIDx). The results of other methods on CheXpert and RSNA are from original paper Wang et al. (2022). The methods including UniMedI shown in the table are pre-trained on MIMIC-CXR dataset, which achieves a fair comparison. As for the state-of-the-art method, MGCA, we mainly compare the performance with the MGCA (ViT-B/16) which employs the ViT as the visual encoder. It is obvious that our method shows the best performance in the three 2D medical image classification for the different training data ratio (1%, 10%, 100%), outperforming the state-of-the-art MGCA (ViT-B/16) by a large margin. Specifically, our method outperforms MGCA with ViT-B/16 backbone with +0.6%, +0.6%, +0.8% AUROC on CheXpert dataset, +0.9%, +0.5%, +0.7% AUROC on RSNA dataset and +5.5%, +7.6%, +2.3% ACC on COVIDx dataset under the 1%, 10%, 100% training ratio respectively. The significant improvement indicates the data efficiency and effectiveness of our method.

### v2

@@ 3.2 Creating Pseudo-Pairs via Attentive Slice Selection
There are $L$ layers with $l$ as the layer index and $H$ heads with $h$ as the head index. $\mathbf{q}_{lh}(\left[CLS\right])\in\mathbb{R}^{C}$ denotes the query embedding of the $\left[CLS\right]$ token at layer $l$ and head $h$. $\mathbf{K}_{lh}(\left[Patch\right])\in\mathbb{R}^{P\times C}$ denotes the key embedding of the all image patch tokens at layer $l$ and head $h$. $C$ is the dimension of the query and key embedding.
After obtaining the attention value for each token, which corresponds to a block of voxels in a slice of the original 3D image, the score for each slice is calculated by averaging values of all tokens in the slice. Therefore, the relevance score $s_{i}$ for the $i$-th slice is
$s_{i}=\frac{1}{N}\sum_{j=1}^{N}\mathbf{v}_{ij}.$ (2)
Here, $\mathbf{v}_{ij}$ is the attention value for the $j$-th token in $i$-th slice that calculated via Eq 1. $N$ represents the number of tokens in a slice. Hence, the relevance score of each slice to the report is obtained via the above average aggregating, and top $k$ slices are selected to establish a connection between 3D and 2D images, allowing learning a shared semantic feature space.
Figure 4: Illustration for the attentive slice selection strategy that creates pseudo-pairs for 2D and 3D medical images.
3.3 Enabling Language Supervisions via VL Contrastive Learning
Generally, UniMedI employs the VL contrastive learning loss in CLIP [29] to introduce language supervisions in reports for learning representations of multi-modal medical images, including both 2D and 3D ones, and build the alignment between language and images. For 2D images (i.e., X-rays), we directly use $T_{2D}$ and $E_{v}$ for feature extraction, obtaining the $\left[CLS\right]$ token that contains global image information, and then align it with the $\left[CLS\right]$ token generated by the language encoder $E_{l}$ that contains global report information. For 3D images (i.e., CT scans), we first select 2D slices related to the report from the 3D volume via the strategy described in Section 3.2. Then, 2D slices and the original 3D volume are tokenized by $T_{2D}$ and $T_{3D}$, respectively. All tokens are concatenated and processed together by $E_{v}$. Finally, same as 2D images, the contrastive learning loss $L_{vl}$ is applied to image $\left[CLS\right]$ tokens and language $\left[CLS\right]$ tokens.
A highlight of UniMedI is the synergistic effect of the attentive slice selection strategy and the VL contrastive learning. On the one hand, VL contrastive learning enables language supervisions, which is directly applied to the visual $\left[CLS\right]$ token, and thus this token contains important information in reports. This is the reason why attention weights of the visual $\left[CLS\right]$ token are utilized as the basis for 2D slice selection. Only in this way can chosen 2D slices carry the supervision information from the report and, together with the 3D features, construct a joint feature space. Here, the key insight is language is used as the coordinate for learning a common semantic space. On the other hand, the attentive slice selection make 2D and 3D feature space more integrated, even without paired data. Thus, they can benefit more from each other, and this common space can amplify similar information between medical images and reports, as they are all about an individual’s health status. In this way, the alignment between images and reports is impelled. In summary, this two designs bring together representations of both 2D and 3D medical images, and make them close to the report representation space at the same time, achieving the effect of one plus one greater than two on building the common semantic space.
3.4 Enhancing Cross-Dimension Interactions via Self-distillation
As medical images themselves are multi-modal and also share some common information, to make the most of these information, we introduce an auxiliary task in which tokens of 2D and 3D images can communicate with each other. Therefore, cross-dimension interactions are reinforced in this way, further enhancing the integration of 2D and 3D images.
We adopt a simple and straightforward auxiliary task design, which is masking and recovering, and use the self-distillation method to complete the task [38, 42], due to its simplicity and effectiveness. Specifically, during the training process, we maintain an online student network $E_{v}$, and a teacher network $\overline{E}_{v}$, which is updated by exponential moving averaged (EMA) over $E_{v}$. The teacher network provides the distillation target as follows. We mask a large proportion of both 2D and 3D image tokens for inputs to the student network, while keeping them complete for inputs to the teacher network. Then, the student network is required to output similar features as the teacher network, even if a significant amount of information is missed in its inputs.
The head and loss of DINO [5] is used for implementing self-distillation here. This loss is applied to both the global $\left[CLS\right]$ tokens (i.e., image contrastive learning loss $L_{icl}$) and all local patch tokens (i.e., patch contrastive learning loss $L_{pcl}$), thereby enabling interactions at different levels of granularity to enhance feature robustness. Besides, since language guidance is the key to realize the common semantic space, the mask strategy here is also based on attention weights, similar to the slice selection method. Tokens with highest attention weights are retained.
Since a large ratio of inputs are masked, this non-trivial auxiliary task forces the model to dig more cross-dimension information, implicitly requiring the model to recover 2D images from 3D ones and vice versa. Therefore, 2D and 3D image tokens can communicate with each other, and cross-dimension interactions are enhanced. Besides, the language-guided masking strategy also makes the interactions more meaningful.
4 Experiments
4.1 Data and Settings in the Pre-Training Stage
4.1.1 Dataset
We pre-train our UniMedI framework on the JPG version of 2D X-rays dataset MIMIC-CXR 2.0.0 [17] and 3D CT scans dataset BIMCV [15]. Specially, for the preprocessing of 2D datasets, we eliminate all side-view images to align with downstream tasks that solely employ front-view images. Additionally, we discard any brief reports that contain fewer than three sentences for 2D and 3D dataset to preserve the integrity of our dataset.
4.1.2 Implementation Details
We utilize ViT-B/16 [10] as the vision encoder to extract representations in the common feature space for 2D and 3D visual data. For text feature encoding, we use BioClinicalBERT [1] as the text encoder to obtain the report embeddings. The vision encoder and text encoder are universal among 2D X-rays and 3D CT scans data. It is worth noting that the patch embed module of vision encoder has different operations for 2D X-rays and 3D CT scans. In general, the image size of 2D images is $224\times 224$ and the volume size of 3D volumes is $128\times 128\times 32$. We pre-train our UniMedI framework 50 epochs on 8 Tesla V100 GPUs with batch size of 144. The optimizer is AdamW [23] with learning rate of $2e^{-5}$ and weight decay of 0.05, where the learning rate follows a linear warmup with cosine annealing scheduler [22]. We initialize learning rate as $1e^{-8}$ and warmup epoch as 20.
4.2 Downstream Tasks for Evaluation
4.2.1 Medical Image Classification
We conduct medical image classification on both 2D and 3D datasets. First, we explore three representative 2D datasets: (1) CheXpert [16], which contains 191,229 frontal-view chest radiographs. The task is to classify each image into 5 individual binary labels: atelectasis, cardiomegaly, consolidation, edema, and pleural effusion. Following [40, 14], we hold out the expert-labeled validation set as test data and randomly select 5,000 radiographs from training data for validation. (2) RSNA Pneumonia [31]. We use the stage 2 version, which contains around 29,700 frontal view chest radiographs. The task is a binary classification, i.e., classifying each chest image into normal or pneumothorax positive. Following [14], we manually split the dataset into training, validation, and test set with 70%/15%/15% ratio. (3) COVIDx [35], which contains over 30k CXR images from a multinational cohort of over 16,600 patients. This dataset contains 16,490 positive COVID-19 images from over 2,800 patients. We use the latest version 6 of this dataset. The task is a three-class classification, i.e., classifying each radiograph into COVID-19, non-COVID pneumonia or normal. We use the original validation dataset as test data and manually split 10% of original training set for validation.
We conduct medical volume classification on two representative datasets: CC-CCII [39] and LUNA 16 [30]. (1) CC-CCII, we use a clean version named Clean-CC-CCII, which consists of 340,190 slices of 3,993 scans from 2,698 patients. The task is to classify each volume into three categories: novel coronavirus pneumonia (NCP), common pneumonia (CP), and normal (Normal). (2) LUNA 16 [30], which is established from LIDC-IDRI [2]. The dataset contains 888 CT scans with annotations which removes CT scans with slice thickness greater than 3mm of LIDC-IDRI database. The task is a binary classification, i.e., classifying each CT volume into pulmonary nodule or normal.
Linear classification is employed to assess the representational capacity of our universal vision-language pre-training model. This model involves keeping the pre-trained ViT vision encoder fixed and solely training a linear classification head initialized randomly for the classification task with varying amounts of training data on each dataset. In addition to this, we perform full model fine-tuning to assess the effectiveness of UniMedI on 3D CT scans.
4.2.2 Medical Image Semantic Segmentation
To assess the effectiveness of UniMedI for medical semantic segmentation, we carry out experiments across a range of representative 2D and 3D datasets. First, we conduct experiments on a representative 2D dataset. RSNA Pneumonia [31], contains 29700 frontal view radiograph. The task is to predict bounding boxes indicating evidence of pneumonia. We randomly split the original training set into 16,010/5,337/5,337 for training/validation/testing. We convert object detection ground truths into masks for semantic segmentation. (2) BCV [18], which consists of 50 CT scans. We partition the training and testing sets following [36].
4.3 Compared Methods
To evaluate the effectiveness of UniMedI, we conduct comparative analyses with various medical VLP methods (ConVIRT [40], GLoRIA [14], MGCA [34], LOVT [24], PRIOR [7]) . However, it is important to note that these methods, different from ours, are tailored specifically for X-ray and their corresponding medical reports. In addition to this, we have also compared our method with several 2D and 3D co-learning methods [36, 25]. UniMiSS not only learns 2D and 3D representations but also concurrently learn all 2D sections derived from 3D volumes, along with all 2D X-ray data. Joint directly learn all 2D sections derived from 3D volumes, along with all 2D X-ray data through contrastive learning.
Table 1: 2D Linear classification results on CheXpert, RSNA and COVIDx with 1%, 10%, 100% training data. Area under ROC curve (AUROC [%]) are reported for CheXpert and RSNA dataset, and accuracy (ACC [%]) is reported for COVIDx dataset. The best results are highlighted in boldface.
@@ 4.3 Compared Methods
DSVE [11] 50.1 51.0 51.5 49.7 52.1 57.8 - - -
VSE++ [12] 50.3 51.2 52.4 49.4 57.2 57.9 - - -
GLoRIA [14] 86.6 87.8 88.1 86.1 88.0 88.6 67.3 77.8 89.0
@@
Caption-Transformer [8] 77.2 82.6 83.9 - - - - - -
Caption-LSTM [37] 85.2 85.3 86.2 - - - - - -
Contrastive-Binary [32] 84.5 85.6 85.8 - - - - - -
ConVIRT [40] 85.9 86.8 87.3 77.4 80.1 81.3 72.5 82.5 92.0
GLoRIA-MIMIC [14] 87.1 88.7 88.0 87.0 89.4 90.2 66.5 80.5 88.8
LOVT [24] 85.1 88.1 88.3 85.5 86.5 89.3 - - -
PIROR [7] 86.2 88.3 88.6 85.7 87.1 89.2 - - -
MGCA (ResNet-50) [34] 87.6 88.0 88.2 88.6 89.1 89.9 72.0 83.5 90.5
MGCA (ViT-B/16) [34] 88.8 89.1 89.7 89.1 89.9 90.8 74.8 84.8 92.3
UniMedI (ViT-B/16) 89.4 89.7 90.5 90.0 90.4 91.5 80.3 92.4 94.6
Table 2: 3D Linear classification results on CC-CCII with 1%, 10%, 100% training data. Accuracy are reported for the dataset. $\ast$ denotes the input size $16\times 96\times 96$. Others is $32\times 128\times 128$. The best results are highlighted in boldface.
@@
Table 3: 3D Fine-tuning results on CC-CCII, RICORD with full training data. ACC [%] is reported for CC-CCII and AUC [%] is reported for LUNA2016-v2. The best results are highlighted in boldface.
@@
4.4 Experiment Results
4.4.1 Results on 2D Medical Image Classification
@@
Table 1 reports the results of linear classification on three 2D medical image classification datasets (CheXpert, RSNA and COVIDx). The results of other methods on CheXpert and RSNA are from original paper. The methods including UniMedI shown in the table are pre-trained on MIMIC-CXR dataset, which achieves a fair comparison. As for the state-of-the-art method, MGCA, we mainly compare the performance with the MGCA (ViT-B/16) which employs the ViT as the visual encoder. It is obvious that our method shows the best performance in the three 2D medical image classification for the different training data ratio (1%, 10%, 100%), outperforming the state-of-the-art MGCA (ViT-B/16) by a large margin. Specifically, our method outperforms MGCA with ViT-B/16 backbone with +0.6%, +0.6%, +0.8% AUROC on CheXpert dataset, +0.9%, +0.5%, +0.7% AUROC on RSNA dataset and +5.5%, +7.6%, +2.3% ACC on COVIDx dataset under the 1%, 10%, 100% training ratio respectively. The significant improvement indicates the data efficiency and effectiveness of our method.
4.4.2 Results on 3D Medical Volume Classification